OpenStudy (anonymous):

Triple integration with spherical coordinates: ∫∫∫E e^(sqrt[x^2+y^2+z^2]); E is bounded by the xz-plane and the hemispheres: y=sqrt[9-x^2-z^2] and y=sqrt[16-x^2-z^2]

OpenStudy (anonymous):

We first need to figure out what E looks like. The equations \[y=\sqrt{9-x ^{2}-z ^{2}} \] and \[y=\sqrt{16-x ^{2}-z ^{2}}\] describe the right halves of spheres with radii 3 and 4, respectively. The xz-plane gives us the left bound of E. So E is the region between the spheres of radii 3 and 4 and to the right of the xz-plane. Using spherical coordinates, the bounds on E are \[0\le \theta \le \pi\] \[0\le \phi \le \pi\] \[3\le \rho \le4\] and knowing that \[\sqrt{x ^{2}+y ^{2}+z ^{2}}=\rho\] We write \[\int\limits_{}^{}\int\limits_{}^{}\int\limits_{}^{}E dV=\int\limits_{3}^{4}\int\limits_{0}^{\pi}\int\limits_{0}^{\pi}e ^{\rho}*\rho ^{2}\sin \phi d \phi d \theta d \rho \] \[=\int\limits_{3}^{4}\int\limits_{0}^{\pi}2\rho ^{2}e ^{\rho}d \theta d \rho\] \[=\int\limits_{3}^{4}2\pi \rho ^{2}e ^{\rho}d \rho\] At this point you can just integrate by parts....